# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 3.5 Diamagnetic Atomic EDMs
The Schiff moment component is sensitive to the neutron and proton EDM and to $PT$-odd nucleon-nucleon interactions. These have been classified in phenomenological one-meson-exchange models Barton:1961eg ; Towner:1994qe ; Engel:1999np , and in chiral EFT and pionless EFT Maekawa:2011vs ; deVries:2012ab ; Song:2012yh ; deVries:2020iea ; Yang:2020ges . For chiral-symmetry breaking operators, such as the qCEDM and the four-quark operators, $\bar{g}_{0,1}$ give the leading contribution to the $PT$-odd nucleon-nucleon potential Maekawa:2011vs ; deVries:2012ab , even though short-range operators in the ${}^{3}P_{0}$-${}^{1}S_{0}$ channel are required for renormalization deVries:2020loy . For chiral-invariant operators, such as the Weinberg three-gluon operator, pion-nucleon couplings and short-range isospin-invariant $NN$ couplings contribute at the same order Maekawa:2011vs ; deVries:2012ab . Since little can be said about the size of the $NN$ CP-odd short-range interactions induced by WET operators, we focus here on the contributions to the Schiff moment from $\bar{g}_{0,1}$. With progress from Lattice QCD and nuclear theory, the Schiff moment expressions could in the future be extended to include additional operators.
@@
The prefactor $A_{\rm Schiff}$ account for the Schiff screening, while the coefficients $\alpha_{n,p}$ and $a_{0,1,2}$ contain the information on the nuclear structure of the systems of interest. We report them in Table 3. We use the results of Ref. Hubert:2022pnl for the screening factors of Hg and Xe, and Ref. Prasannaa:2020cjx for 225Ra. In the case of 199Hg, $\alpha_{n,p}$ are known with good accuracy Dmitriev:2003sc , while in the case of Xe we use the calculation of $\alpha_{n}$ from Yanase:2022atk . The contributions of the CP-odd pion-nucleon couplings are affected by substantial uncertainties. Here we take the ranges suggested in Refs. Engel:2013lsa ; Chupp:2017rkp , which considers results from various nuclear structure calculations.
@@
$\left.\mathcal{R}_{\rm T}\right|_{\rm Hg}=-\left(2.8\pm 0.6\right)\cdot 10^{-20}e\,{\rm cm}\qquad\left.\mathcal{R}_{\rm T}\right|_{\rm Xe}=\left(0.520\pm 0.049\right)\cdot 10^{-20}e\,{\rm cm}.$ (76)
@@ 4 Renormalization group running and Threshold Effects
  • One-loop threshold effects at $\mu=m_{c}$ and $\mu=m_{b}$.
@@ 4.1 Leading-log 1-loop RG running
Figure 1: Representative diagrams contributing to the one-loop anomalous dimension, which induces LL contributions to EDMs. Plain and double lines denote light ($e$ and $u$, $d$, $s$ quarks) and heavy fermions ($\mu$, $\tau$ leptons and $c$, $b$ quarks), respectively. Wavy and wiggly lines denote photons and gluons. Insertions of WET operators are denoted by squares, while SM vertices by dots. Diagrams $(a)$ and $(b)$ contribute to the mixing of heavy-light four-fermion operators onto photon and gluon dipole operators. Four-quark operators of the type $\mathcal{O}_{uu}^{S1,RR}$, $\mathcal{O}_{uu}^{S8,RR}$, $\mathcal{O}_{dd}^{S8,RR}$, $\mathcal{O}_{dd}^{S8,RR}$, $\mathcal{O}_{uddu}^{S1,RR}$ and $\mathcal{O}_{uddu}^{S8,RR}$, with two heavy and two light quarks, semileptonic tensor operators $\mathcal{O}^{T,RR}_{eu}$ $\mathcal{O}^{T,RR}_{ed}$ with heavy quarks or heavy leptons, and the leptonic scalar operator $\mathcal{O}^{S,RR}_{ee}$ with two heavy and two light leptons generate LL EDMs via this path. Diagram $(c)$ denotes the mixing of dipole operators. Diagram $(d)$ denotes the mixing of the semileptonic scalar operators $\mathcal{O}^{S,RR}_{eu}$ and $\mathcal{O}^{S,RR}_{ed}$ onto tensor operators. Combined with the running of the tensor into dipoles, these induce $1/(4\pi)^{4}\log^{2}$ corrections to the electron EDM.
The complete 1-loop anomalous dimension matrices (ADMs) in WET including $\Delta F=0$ operators have been calculated in Jenkins:2017dyc . In what follows, we identify all relevant terms in the ADMs and provide the naive solutions for the corresponding RGEs. Through diagrams such as diagram $(a)$ and $(c)$ in Figure 1, the QCD dipole and four-quark scalar operators mix onto the photon dipole. Neglecting the running of the electromagnetic and strong couplings, and the running of the Wilson coefficients, we can denote the solution of the RGEs as
$\big{[}{C}_{f\gamma}\big{]}_{ii}(\mu_{low})$ $={8eg_{s}q_{i}C_{F}}\big{[}{C}_{fG}\big{]}_{ii}L_{\mu}\,,$ (80)
$\big{[}{C}_{f\gamma}\big{]}_{ii}(\mu_{low})$ $={2eq_{j}m_{j}}\big{[}{C}_{ff}^{S1,RR}\big{]}_{ijji}L_{\mu}\,,$
$\big{[}{C}_{f\gamma}\big{]}_{ii}(\mu_{low})$ $={2eq_{j}m_{j}C_{F}}\big{[}{C}_{ff}^{S8,RR}\big{]}_{ijji}L_{\mu}\,,$
$\big{[}{C}_{f\gamma}\big{]}_{ii}(\mu_{low})$ $=-8eq_{e}m_{j}\big{[}{C}_{ef}^{T,RR}\big{]}_{jjii}L_{\mu}\,,$
$\big{[}{C}_{u\gamma}\big{]}_{ii}(\mu_{low})$ $={eq_{j}m_{j}}\big{[}{C}_{uddu}^{S1,RR}\big{]}_{ijji}L_{\mu}\,,$
$\big{[}{C}_{u\gamma}\big{]}_{ii}(\mu_{low})$ $={eq_{j}m_{j}C_{F}}\big{[}{C}_{uddu}^{S8,RR}\big{]}_{ijji}L_{\mu}.$
@@
$\big{[}{C}_{aG}\big{]}_{ii}$ $={2g_{s}m_{j}}\big{[}{C}_{aa}^{S1,RR}\big{]}_{ijji}L_{\mu}\,,$ (82)
$\big{[}{C}_{aG}\big{]}_{ii}$ $={2g_{s}m_{j}}\left(C_{F}-{C_{A}\over 2}\right)\big{[}{C}_{aa}^{S8,RR}\big{]}_{ijji}L_{\mu}\,,$
@@
The mixing of $\big{[}{C}_{a\gamma}\big{]}$ onto $\big{[}{C}_{aG}\big{]}$ follows the exact same relation as the first equation in (80). Also, the mixing of $\big{[}{C}_{uddu}^{S1(S8),RR}\big{]}$ onto $\big{[}{C}_{dG}\big{]}$ follows the same relations as the last two equations in (82). Eqs. (80) and (82) show that all scalar operators with two heavy and two light quarks induce light quark EDMs and CEDMs at one loop, leading, as we will see, to strong constraints from neutron and mercury EDM experiments. For the electron EDM, the mixing of the scalar and tensor operators onto the dipole operator is governed by pure QED
@@
Finally, through diagrams such as diagram $(d)$ in Fig. 1, the leptonic operators $\big{[}{C}_{ee}^{S,RR}\big{]}_{1122}$ and $\big{[}{C}_{ee}^{S,RR}\big{]}_{1133}$ mix with $\big{[}{C}_{e\gamma}^{S,RR}\big{]}_{11}$ via $\big{[}{C}_{ee}^{S,RR}\big{]}_{1221}$ and $\big{[}{C}_{ee}^{S,RR}\big{]}_{1331}$ operators. Similarly, semileptonic scalar operators mix into tensor operators, and scalar operators of the form $\big{[}{C}_{uu}^{S1(8),RR}\big{]}_{iijj}$, $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{iijj}$ and $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{iijj}$ mix onto $\big{[}{C}_{uu}^{S1(8),RR}\big{]}_{ijji}$, $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{ijji}$ and $\big{[}{C}_{uddu}^{S1(8),RR}\big{]}_{ijji}$, respectively. In this way, these operators induce LL contribution to the electron and quark dipole operators, which however start at two loops and go as $L_{\mu}^{2}$.
@@
This exhaust all the operators denoted as LL in Table 1.
@@ 4.2.1 Matching to Weinberg operators
The 1-loop Feynman diagram governing this mixing is given in Fig. 2, diagram $(a)$. Similarly, we can integrate out the charm quark leading to
@@ 4.2.2 Matching to dipole operators
The vector operators with left-right chiral currents match onto dipole operators at 1-loop Alioli:2017ces . For chromo-dipoles, in the JMS basis, we find
@@
similarly, using the last two equations one can obtain matching for $\big{[}{C}_{uG}\big{]}_{jj}$ and $\big{[}{C}_{dG}\big{]}_{ii}$ For the photon dipoles, we have
@@
In the above relations, the scale $\mu$ can be set to the appropriate matching scale of either $m_{c}$ or $m_{b}$. Finally, for the leptonic dipole, at $\mu=m_{\tau}$, we obtain Aebischer:2021uvt
@@ 4.3 Two-step operator mixing
Figure 3: Representative diagrams that induce NLL contributions to EDMs. Diagrams $(a)$ and $(b)$ denote the LL mixing of scalar four-fermion operators with four heavy quarks onto heavy flavor dipoles, while diagram $(c)$ the running of heavy flavor dipoles into themselves. At the heavy flavor threshold, the gluonic dipole is integrated out, inducing $\mathcal{O}((4\pi)^{-4}\log)$ contributions to $C_{\tilde{G}}$. Diagrams $(d)$ and $(e)$ contribute to the two-loop mixing of heavy-light vector operators onto dipoles.
We next consider possible NLL effects, which give rise to the dominant EDM contributions for those operators that do not contribute at LL. As shown in (89) and (90) the QCD dipoles $\big{[}{C}_{dG}\big{]}_{33}$ and $\big{[}{C}_{uG}\big{]}_{22}$ can mix with the three-gluon operator $\big{[}{C}_{\widetilde{G}}\big{]}$ due to 1-loop threshold corrections at the bottom and charm thresholds, respectively. As illustrated in diagram $(b)$ in Fig. 3, the scalar operators $\big{[}{C}_{uu}^{S1,RR}\big{]}_{2222}$ and $\big{[}{C}_{uu}^{S8,RR}\big{]}_{2222}$ mix with $\big{[}{C}_{uG}\big{]}_{22}$ at 1-loop level. Therefore, $\big{[}{C}_{uu}^{S1,RR}\big{]}_{2222}$ and $\big{[}{C}_{uu}^{S8,RR}\big{]}_{2222}$ induce a correction to the three gluon operator of the form
@@
$L_{\mu_{c}}=\frac{1}{16\pi^{2}}\log\left(\frac{\mu_{ew}}{m_{c}}\right),\qquad L_{\mu_{b}}=\frac{1}{16\pi^{2}}\log\left(\frac{\mu_{ew}}{m_{b}}\right),$ (96)
@@
The operators $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{2233}$ have one-loop QCD and QED mixing with $\big{[}{C}_{uddu}^{S1(8),RR}\big{]}_{2332}$, and thus generate contributions to the $b$ and $c$ QED and QCD dipole analogous to Eqs. (87) and (88). These generate corrections to $\big{[}{C}_{\widetilde{G}}\big{]}$ that scale as $L^{2}_{\mu}/(16\pi^{2})$.
@@
We have seen that four-quark scalar operators in the $(LR)(LR)$ class contribute to EDM at LL, in the case of operators with four light quarks or two heavy and two light quarks, or NLL, for operators with four heavy quarks. We next analyze vector operators in the $(LL)(RR)$ class. For heavy-light operators, the first contribution arises at NLL. At this order, we need to consider the two-loop running between $\mu_{ew}$ and $m_{b}$ or $m_{c}$, and the one-loop matching contribution discussed in Sec. 4.2.2. Both the anomalous dimension and the matching coefficient are separately scheme-dependent, and the scheme dependence cancels between the two Ciuchini:1993ks . The two-loop anomalous dimension can be obtained by generalizing the results of Misiak and Cho Cho:1993zb for $\Delta F=1$ operators in the standard basis to $\Delta F=0$ four-quark and leptonic operators for mixing between the vector left-right and dipole operators in the JMS basis. A typical 2-loop Feynman diagram governing this mixing is given by diagram $(d)$ in Fig. 3. The effective Hamiltonian for $\Delta F=1$ was used to be Cho:1993zb
$\mathcal{H}_{\rm eff}^{\Delta F=1}=V_{tb}V_{ts}^{*}\frac{G_{F}}{\sqrt{2}}\sum_{i\in 7,8,9,10}C_{i}Q_{i}.$ (98)
@@
Here the superscript $u$ indicates that the $u$-quarks are closed in diagram $(d)$ in Fig. 3. Since the JMS basis uses the unit normalization for the operators, using (103), first go to an intermediate basis and obtain the ADM for it
@@
Here, the relation between ${\hat{\gamma}}^{\prime u}$ and ${\hat{\gamma}}_{S}^{u}$ is
@@
The matrix $R(\mu)$ is the transformation matrix defined by $C^{\prime}(\mu)=R(\mu)C_{S}(\mu)$. The RGE fo $R(\mu)$ is dictated by the RGEs of $g_{s}(\mu)$ and $m_{b}(\mu)$
@@
Here we have used $16\pi^{2}\mu({dm_{b}/d\mu})=-6C_{F}g_{s}^{2}m_{b}$ and $316\pi^{2}\mu({dg_{s}/d\mu})=-23g_{s}^{3}/$ 3 for five flavors. The correspondence between the $Q^{\prime}$ and the JMS bases is given by
@@
In this case, the operators ${C_{\!\underset{prrp}{uddu}}^{{\tiny V1,LR}}}$ and ${C_{\!\underset{prrp}{uddu}}^{{\tiny V8,LR}}}$ contribute to EDMs, here the index $p$ can be equal or unequal to $r$. At 2-loop order, these two operators can mix with $\{{C_{\!\underset{pp}{u\gamma}}},{C_{\!\underset{pp}{uG}}}\}$, if one closes the $d$ quark loop, or $\{{C_{\!\underset{rr}{d\gamma}}},{C_{\!\underset{rr}{dG}}}\}$, by closing the $u$ quark loop. Using the color identity $T^{A}_{ij}A^{A}_{kl}={1\over 2}\left(\delta_{il}\delta_{jk}-{1\over 3}\delta_{ij}\delta_{kl}\right)$, the tree-level transformation matrix between the intermediate basis with properly adjusted down-quark flavors to $\Delta F=0$ and the JMS basis $\vec{\mathcal{C}}_{u}=\{{C_{\!\underset{rr}{d\gamma}}},{C_{\!\underset{rr}{dG}}},$ ${C_{\!\underset{prrp}{uddu}}^{{\tiny V1,LR}}},{C_{\!\underset{prrp}{uddu}}^{{\tiny V8,LR}}}\}$ is found to be
@@
Using (101), (106) and (108) in (105), we obtain the ADMs for $\vec{\mathcal{C}}_{u}$ when the $u$-quark loop is closed [^3] Note that in the original result of Misiak and Cho the charm-quark loop was closed.
$\gamma_{\text{JMS}}^{u}=\begin{pmatrix}8g_{s}^{2}\over 3&-32eg_{s}\over 9&4eg_{s}^{2}m_{u}\over 9\pi^{2}&43eg_{s}^{2}m_{u}\over 27\pi^{2}\\ 0&-19g_{s}^{2}\over 3&g_{s}^{3}m_{u}\over 6\pi^{2}&-11g_{s}^{3}m_{u}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (109)
@@
$\gamma_{\text{JMS}}^{d}=\begin{pmatrix}8g_{s}^{2}\over 3&-32eg_{s}\over 9&eg_{s}^{2}m_{d}\over 9\pi^{2}&-23eg_{s}^{2}m_{d}\over 27\pi^{2}\\ 0&-19g_{s}^{2}\over 3&g_{s}^{3}m_{d}\over 6\pi^{2}&-11g_{s}^{3}m_{d}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (110)

### v2

@@ 3.5 Diamagnetic Atomic EDMs
The Schiff moment component is sensitive to the neutron and proton EDM and to $PT$-odd nucleon-nucleon interactions. These have been classified in phenomenological one-meson-exchange models Barton:1961eg ; Towner:1994qe ; Engel:1999np , and in chiral EFT and pionless EFT Maekawa:2011vs ; deVries:2012ab ; Song:2012yh ; deVries:2020iea ; Yang:2020ges . For chiral-symmetry breaking operators, such as the qCEDM and the four-quark operators, $\bar{g}_{0,1}$ give the leading contribution to the $PT$-odd nucleon-nucleon potential Maekawa:2011vs ; deVries:2012ab , even though short-range operators in the ${}^{3}P_{0}$-${}^{1}S_{0}$ channel are required for renormalization deVries:2020loy . For chiral-invariant operators, such as the Weinberg three-gluon operator, $\pi N$ couplings and short-range isospin-invariant $NN$ couplings contribute at the same order Maekawa:2011vs ; deVries:2012ab . Since little can be said about the size of the $NN$ CP-odd short-range interactions induced by WET operators, we focus here on the contributions to the Schiff moment from $\bar{g}_{0,1}$. With progress from Lattice QCD and nuclear theory, the Schiff moment expressions could in the future be extended to include additional operators.
@@
The prefactor $A_{\rm Schiff}$ accounts for the Schiff screening, while the coefficients $\alpha_{n,p}$ and $a_{0,1,2}$ contain the information on the nuclear structure of the systems of interest. We report them in Table 3. We use the results of Ref. Hubert:2022pnl for the screening factors of Hg and Xe, and Ref. Prasannaa:2020cjx for 225Ra. In the case of 199Hg, $\alpha_{n,p}$ are known with good accuracy Dmitriev:2003sc , while in the case of Xe we use the calculation of $\alpha_{n}$ from Yanase:2022atk . The contributions of the CP-odd $\pi N$ couplings are affected by substantial uncertainties. Here we take the ranges suggested in Refs. Engel:2013lsa ; Chupp:2017rkp , which considers results from various nuclear structure calculations.
@@
$\left.\mathcal{R}_{\rm T}\right|_{\rm Hg}=-\left(2.8\pm 0.6\right)\cdot 10^{-20}e\,{\rm cm}\qquad\left.\mathcal{R}_{\rm T}\right|_{\rm Xe}=\left(0.520\pm 0.049\right)\cdot 10^{-20}e\,{\rm cm},$ (76)
@@ 4 Renormalization group running and threshold effects
  • 1-loop threshold effects at $\mu=m_{c}$ and $\mu=m_{b}$.
@@ 4.1 Leading-log 1-loop RG running
Figure 1: Representative diagrams contributing to the 1-loop anomalous dimension, which induces LL contributions to EDMs. Plain and double lines denote light ($e$ and $u$, $d$, $s$ quarks) and heavy fermions ($\mu$, $\tau$ leptons and $c$, $b$ quarks), respectively. Wavy and wiggly lines denote photons and gluons. Insertions of WET operators are denoted by squares, while SM vertices by dots. Diagrams $(a)$ and $(b)$ contribute to the mixing of heavy-light four-fermion operators onto photon and gluon dipole operators. Four-quark operators of the type $\mathcal{O}_{uu}^{S1,RR}$, $\mathcal{O}_{uu}^{S8,RR}$, $\mathcal{O}_{dd}^{S8,RR}$, $\mathcal{O}_{dd}^{S8,RR}$, $\mathcal{O}_{uddu}^{S1,RR}$ and $\mathcal{O}_{uddu}^{S8,RR}$, with two heavy and two light quarks, semileptonic tensor operators $\mathcal{O}^{T,RR}_{eu}$ $\mathcal{O}^{T,RR}_{ed}$ with heavy quarks or heavy leptons, and the leptonic scalar operator $\mathcal{O}^{S,RR}_{ee}$ with two heavy and two light leptons generate LL EDMs via this path. Diagram $(c)$ denotes the mixing of dipole operators. Diagram $(d)$ denotes the mixing of the semileptonic scalar operators $\mathcal{O}^{S,RR}_{eu}$ and $\mathcal{O}^{S,RR}_{ed}$ onto tensor operators. Combined with the running of the tensor into dipoles, these induce corrections to the electron EDM starting at $\mathcal{O}(L_{\mu}^{2})$.
The complete 1-loop ADMs in WET including $\Delta F=0$ operators have been calculated in Jenkins:2017dyc . In what follows, we identify all relevant terms in the ADMs and provide the naive solutions for the corresponding RGEs. Through diagrams such as diagram $(a)$ and $(c)$ in Figure 1, the QCD dipole and four-quark scalar operators mix onto the photon dipole. Neglecting the running of the electromagnetic and strong couplings, and the running of the Wilson coefficients, we can denote the solution of the RGEs as
$\big{[}{C}_{f\gamma}\big{]}_{ii}$ $={8eg_{s}q_{i}C_{F}}\big{[}{C}_{fG}\big{]}_{ii}L_{\mu}\,,$ (80)
$\big{[}{C}_{f\gamma}\big{]}_{ii}$ $={2eq_{j}m_{j}}\big{[}{C}_{ff}^{S1,RR}\big{]}_{ijji}L_{\mu}\,,$
$\big{[}{C}_{f\gamma}\big{]}_{ii}$ $={2eq_{j}m_{j}C_{F}}\big{[}{C}_{ff}^{S8,RR}\big{]}_{ijji}L_{\mu}\,,$
$\big{[}{C}_{f\gamma}\big{]}_{ii}$ $=-8eq_{e}m_{j}\big{[}{C}_{ef}^{T,RR}\big{]}_{jjii}L_{\mu}\,,$
$\big{[}{C}_{u\gamma}\big{]}_{ii}$ $={eq_{j}m_{j}}\big{[}{C}_{uddu}^{S1,RR}\big{]}_{ijji}L_{\mu}\,,$
$\big{[}{C}_{u\gamma}\big{]}_{ii}$ $={eq_{j}m_{j}C_{F}}\big{[}{C}_{uddu}^{S8,RR}\big{]}_{ijji}L_{\mu}.$
@@
$\big{[}{C}_{fG}\big{]}_{ii}$ $={2g_{s}m_{j}}\big{[}{C}_{ff}^{S1,RR}\big{]}_{ijji}L_{\mu}\,,$ (82)
$\big{[}{C}_{fG}\big{]}_{ii}$ $={2g_{s}m_{j}}\left(C_{F}-{C_{A}\over 2}\right)\big{[}{C}_{ff}^{S8,RR}\big{]}_{ijji}L_{\mu}\,,$
@@
The mixing of $\big{[}{C}_{f\gamma}\big{]}$ onto $\big{[}{C}_{fG}\big{]}$ follows the exact same relation as the first equation in (80). Also, the mixing of $\big{[}{C}_{uddu}^{S1(S8),RR}\big{]}$ onto $\big{[}{C}_{dG}\big{]}$ follows the same relations as the last two equations in (82). Eqs. (80) and (82) show that all scalar operators with two heavy and two light quarks induce light quark EDMs and CEDMs at one loop, leading, as we will see, to strong constraints from neutron and mercury EDM experiments. For the electron EDM, the mixing of the scalar and tensor operators onto the dipole operator is governed by pure QED
@@
Finally, through diagrams such as diagram $(d)$ in Fig. 1, the leptonic operators $\big{[}{C}_{ee}^{S,RR}\big{]}_{1122}$ and $\big{[}{C}_{ee}^{S,RR}\big{]}_{1133}$ mix with $\big{[}{C}_{e\gamma}^{S,RR}\big{]}_{11}$ via the $\big{[}{C}_{ee}^{S,RR}\big{]}_{1221}$ and $\big{[}{C}_{ee}^{S,RR}\big{]}_{1331}$ operators. Similarly, semileptonic scalar operators mix into tensor operators, and scalar operators of the form $\big{[}{C}_{uu}^{S1(8),RR}\big{]}_{iijj}$, $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{iijj}$ and $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{iijj}$ mix onto $\big{[}{C}_{uu}^{S1(8),RR}\big{]}_{ijji}$, $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{ijji}$ and $\big{[}{C}_{uddu}^{S1(8),RR}\big{]}_{ijji}$, respectively. In this way, these operators induce LL contribution to the electron and quark dipole operators, which however start at two loops and go as $L_{\mu}^{2}$.
@@
This exhausts all the operators denoted as LL in Table 1.
@@ 4.2.1 Matching to Weinberg operators
The coefficient of second term is $\approx 7.1\cdot 10^{-4}$. The 1-loop Feynman diagram governing this mixing is given in Fig. 2, diagram $(a)$. Similarly, we can integrate out the charm quark leading to
@@ 4.2.2 Matching to dipole operators
The vector operators with left-right chiral currents match onto dipole operators at one loop Alioli:2017ces . For chromo-dipoles, in the JMS basis, we find
@@
similarly, using the last two equations one can obtain the matching for $\big{[}{C}_{uG}\big{]}_{jj}$ and $\big{[}{C}_{dG}\big{]}_{ii}$ For the photon dipoles, we have
@@
In the above relations, the scale $\mu$ can be set to the appropriate matching scale of either $m_{c}$ or $m_{b}$. Finally, for the leptonic dipoles, at $\mu=m_{\tau}$, we obtain Aebischer:2021uvt
@@ 4.3 Two-step operator mixing
Figure 3: Representative diagrams that induce NLL contributions to EDMs. Diagrams $(a)$ and $(b)$ denote the LL mixing of scalar four-fermion operators with four heavy quarks onto heavy flavor dipoles, while diagram $(c)$ the running of heavy flavor dipoles into themselves. At the heavy flavor threshold, the gluonic dipole is integrated out, inducing $\mathcal{O}((4\pi)^{-2}L_{\mu})$ contributions to $C_{\tilde{G}}$. Diagrams $(d)$ and $(e)$ contribute to the 2-loop mixing of heavy-light vector operators onto dipoles.
We next consider possible NLL effects, which give rise to the dominant EDM contributions for those operators that do not contribute at LL. As shown in (89) and (90) the QCD dipoles $\big{[}{C}_{dG}\big{]}_{33}$ and $\big{[}{C}_{uG}\big{]}_{22}$ can mix with the three-gluon operator $\big{[}{C}_{\widetilde{G}}\big{]}$ due to 1-loop threshold at the bottom and charm thresholds, respectively. As illustrated in diagram $(b)$ in Fig. 3, the scalar operators $\big{[}{C}_{uu}^{S1,RR}\big{]}_{2222}$ and $\big{[}{C}_{uu}^{S8,RR}\big{]}_{2222}$ mix with $\big{[}{C}_{uG}\big{]}_{22}$ at 1-loop level. Therefore, $\big{[}{C}_{uu}^{S1,RR}\big{]}_{2222}$ and $\big{[}{C}_{uu}^{S8,RR}\big{]}_{2222}$ induce a correction to the three gluon operator of the form
@@
$L_{\mu_{c}}=\frac{1}{16\pi^{2}}\log\left(\frac{m_{c}}{\mu_{ew}}\right),\qquad L_{\mu_{b}}=\frac{1}{16\pi^{2}}\log\left(\frac{m_{b}}{\mu_{ew}}\right),$ (96)
@@
The operators $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{2233}$ have 1-loop QCD and QED mixing with $\big{[}{C}_{uddu}^{S1(8),RR}\big{]}_{2332}$, and thus generate contributions to the $b$ and $c$ QED and QCD dipoles analogous to Eqs. (87) and (88). These generate corrections to $\big{[}{C}_{\widetilde{G}}\big{]}$ that scale as $L^{2}_{\mu}/(16\pi^{2})$.
@@
We have seen that four-quark scalar operators in the $(LR)(LR)$ class contribute to EDM at LL, in the case of operators with four light quarks or two heavy and two light quarks, or NLL, for operators with four heavy quarks. We next analyze vector operators in the $(LL)(RR)$ class. For heavy-light operators, the first contribution arises at NLL. At this order, we need to consider the 2-loop running between $\mu_{ew}$ and $m_{b}$ or $m_{c}$, and the 1-loop matching contribution discussed in Sec. 4.2.2. Both the anomalous dimension and the matching coefficient are separately scheme-dependent, and the scheme dependence cancels between the two Ciuchini:1993ks . The 2-loop anomalous dimension can be obtained by generalizing the results of Misiak and Cho Cho:1993zb . They calculated the mixing between $\Delta F=1$ vector left-right and dipole operators in the “standard” basis. Here we generalize their results to to the mixing of $\Delta F=0$ four-fermion and dipole operators in the JMS basis. A typical 2-loop Feynman diagram governing this mixing is given by diagram $(d)$ in Fig. 3. The $\Delta F=1$ effective Lagrangian used in Ref. Cho:1993zb was
$\mathcal{L}_{\rm eff}^{\Delta F=1}=V_{tb}V_{ts}^{*}\frac{G_{F}}{\sqrt{2}}\sum_{i\in 7,8,9,10}C_{i}Q_{i}\,,$ (98)
@@
The superscript $u$ indicates that the $u$-quarks are closed in diagram $(d)$ in Fig. 3. Since the JMS basis uses the unit normalization for the operators, using (103), first we go to an intermediate basis and obtain the ADM for it
@@
The relation between ${\hat{\gamma}}^{\prime u}$ and ${\hat{\gamma}}_{S}^{u}$ is given by
@@
The matrix $R(\mu)$ is the transformation matrix defined by $C^{\prime}(\mu)=R(\mu)C_{S}(\mu)$. The RGE for $R(\mu)$ is dictated by the RGEs of $g_{s}(\mu)$ and $m_{b}(\mu)$
@@
We have used $16\pi^{2}\mu({dm_{b}/d\mu})=-6C_{F}g_{s}^{2}m_{b}$ and $316\pi^{2}\mu({dg_{s}/d\mu})=-23g_{s}^{3}/$ 3 for five flavors. The correspondence between the $Q^{\prime}$ and the JMS bases is given by
@@
In this case, the operators ${C_{\!\underset{prrp}{uddu}}^{{\tiny V1,LR}}}$ and ${C_{\!\underset{prrp}{uddu}}^{{\tiny V8,LR}}}$ contribute to EDMs, here the index $p$ can be equal or unequal to $r$. At 2-loop order, these two operators can mix with $\{{C_{\!\underset{pp}{u\gamma}}},{C_{\!\underset{pp}{uG}}}\}$, if one closes the $d$ quark loop, or $\{{C_{\!\underset{rr}{d\gamma}}},{C_{\!\underset{rr}{dG}}}\}$, by closing the $u$ quark loop. Using the color identity $T^{A}_{ij}T^{A}_{kl}={1\over 2}\left(\delta_{il}\delta_{jk}-{1\over 3}\delta_{ij}\delta_{kl}\right)$, the tree-level transformation matrix between the intermediate basis with properly adjusted down-quark flavors to $\Delta F=0$ and the JMS basis $\vec{\mathcal{C}}_{u}=\{{C_{\!\underset{rr}{d\gamma}}},{C_{\!\underset{rr}{dG}}},$ ${C_{\!\underset{prrp}{uddu}}^{{\tiny V1,LR}}},{C_{\!\underset{prrp}{uddu}}^{{\tiny V8,LR}}}\}$ is found to be
@@
Using (101), (106) and (108) in (105), we obtain  [^3] Note that one needs to properly adjust the value of electric charge for the case of up-type dipoles in contrast to the original result which is valid for the down-quark dipoles. the ADMs for $\vec{\mathcal{C}}_{u}$ when the $u$-quark loop is closed [^4] Note that in the original result of Misiak and Cho the charm-quark loop was closed.
$\gamma_{\text{JMS}}^{u}=\begin{pmatrix}8g_{s}^{2}\over 3&-32eg_{s}\over 9&4eg_{s}^{2}m_{u}\over 9\pi^{2}&43eg_{s}^{2}m_{u}\over 27\pi^{2}\\ -8eg_{s}\over 3&-19g_{s}^{2}\over 3&g_{s}^{3}m_{u}\over 6\pi^{2}&-11g_{s}^{3}m_{u}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (109)
@@
$\gamma_{\text{JMS}}^{d}=\begin{pmatrix}8g_{s}^{2}\over 3&64eg_{s}\over 9&eg_{s}^{2}m_{d}\over 9\pi^{2}&-23eg_{s}^{2}m_{d}\over 27\pi^{2}\\ 16eg_{s}\over 3&-19g_{s}^{2}\over 3&g_{s}^{3}m_{d}\over 6\pi^{2}&-11g_{s}^{3}m_{d}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (110)